OpenStudy (anonymous):

In the function f(x) = 4x - 9 how does an increase of 1 in the independent variable x affect f(x)?

OpenStudy (anonymous):

f(x) will have a decrease of 4x - 4 f(x) will have an increase of 4 f(x) will have an increase of 13 f(x) will have a decrease of 4 are the choices

OpenStudy (anonymous):

instead of putting x, put x+1 into f(x)

OpenStudy (anonymous):

check and see replace x by 2 and compute then replace x by 3 and compute see the difference in the outputs

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

f(x) = 4x - 9 f(x+1) = 4(x+1) -9 expand the right side of the equation. see what happens :)

OpenStudy (anonymous):

f(x+1)=4x+4-9

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

so you have f(x+1) = 4x-5

hartnn (hartnn):

yes, thats correct, so, f(x+1) = (4x-9)+4 = f(x) +4 now, what can you say ?

OpenStudy (anonymous):

i dont know

hartnn (hartnn):

f(x) +4 suggest that f(x) increased by 4, right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the answer is f(x) will have an increase of 4

hartnn (hartnn):

yes., that would be correct :)

OpenStudy (anonymous):

the slope of \(y=mx+b\) is \(m\) meaning an increase in 1 unit in the \(x\) direction gives n increase of \(m\) in the \(y\) direction

OpenStudy (anonymous):

thank you very much

hartnn (hartnn):

welcome ^_^